OpenStudy (anonymous):

How so I find sin(170) degrees without an calculator?

OpenStudy (anonymous):

Sin 170 = sin 10 Sin 10 =Sin (30/3) \[Sin(30)=3\sin(10)-4\sin^3(10)=1/2\] Solve for real solutions of Sin 10

OpenStudy (anonymous):